# Learning Linearity in Audio Consistency Autoencoders via Implicit Regularization

### v1

@@ 1 Introduction
Modern Autoencoders can achieve excellent reconstruction quality at high compression rates at the expense of complex, entangled latent spaces. For applications where input space manipulation is desirable from within the compressed space, recent research has proposed either task-specific adaptors for pre-trained models or redesigning the autoencoder to preserve key structural properties, such as equivariance to spatial transformations [1, 2, 3]). This work follows the latter approach.
@@
This work presents a training methodology for constructing an approximately linear compressed audio representation that enables intuitive manipulation, where simple algebraic operations in the latent space correspond directly to mixing and scaling in the audio domain. The method employs implicit regularization through data augmentation, without modifying the model architecture or objective. The approach is demonstrated with the Music2Latent architecture [4], a Consistency Autoencoder (CAE) that achieves high-quality, single-step reconstruction and a $64\times$ compression rate for $44.1$ kHz audio.
@@ 2.1 Audio processing in the latent space
There is a growing interest in manipulating audio by training task-specific modules which operate directly in the latent space of a pretrained AE, with applications in source separation [5], speech enhancement [6, 7], upsampling and upmixing [8], filtering [9], and generative modelling [10, 11, 12]. Some works retrain AEs for specific objectives, such as source disentanglement [13] or separation via latent masking [14]. In contrast, our work focuses on training an AE to have desirable, task-agnostic structural properties. This enables direct and efficient audio manipulation and can serve as a strong foundation for these downstream applications.
@@ 2.2 Diffusion and Consistency models
Denoising Diffusion Probabilistic Models [15] and score-based models [16] have achieved great success in generative modeling, where the goal is to estimate the underlying data distribution from samples. These models define a forward process that gradually adds noise to data and learn a reverse process to denoise it using a neural network. Sampling typically requires an iterative procedure with many steps to generate a clean sample. Consistency Models [17] accelerate this process by mapping any point along the trajectory defined by the probability flow ordinary differential equation [16] directly to the origin (the clean data point), enabling single-step generation. When trained from scratch, the process is called Consistency Training (CT), in which a ”student” denoiser network ($f_{\theta}$) is trained to match the output of a ”teacher” ($f_{\theta^{-}}$), which itself denoises a less corrupted version of the same data point. In improved Consistency Training (iCT) [18], the teacher is updated with the same parameters $\theta$ as the student but detached from the computational graph ($\theta^{-}\leftarrow\text{stopgrad}(\theta)$).
@@ 2.3 High-Fidelity Audio Autoencoders
In audio, popular AEs include Neural Audio Codecs [19], which compress audio into a set of discrete tokens, and Variational Autoencoders [20], which have seen significant success in generative modeling. However, both frameworks require complex, multi-stage training with adversarial objectives [21] to achieve high-quality reconstructions. Diffusion Autoencoders [22, 23, 24, 25, 10] enable high-fidelity reconstruction by replacing the deterministic decoder with a conditional diffusion model. Notable examples in audio include [10] and Music2Latent [4, 26]. A specific instance of DiffAEs is the CAE[4, 26], where the decoder is a CM, enabling decoding in one step. DiffAEs can be trained with a single diffusion/CT objective [22, 24, 4]. High-quality decoding has been linked to the capacity of sampling details at inference-time, since the decoder acts as a denoiser instead of an upsampler, which also reduces the amount of unnecessary information to be encoded in the latent [23, 25].
@@ 3 Method
We postulate that approximate linearity can be achieved by a simple data augmentation scheme, without the need to change model or loss function. Our approach is in theory model-agnostic and can be applied to any AE, and in this work we apply it on a CAE architecture for audio [4], which offers $64\times$ compression and high-quality reconstruction with a single training objective.
@@ 3.1 Model architecture
Our model (Lin-CAE) follows the Music2Latent architecture [4], illustrated in Figure 2(2(a)). We recall the main components here for clarity, but we refer the reader to [4] for the full details.
CAE Denoiser representation space: We use a complex-valued Short-Time Fourier Transform (STFT) followed by the invertible amplitude scaling ${\operatorname{Amp}}:{\mathbb{C}}\to{\mathbb{C}}=\beta|c|^{\alpha}e^{i\angle(c)}$ (with $\alpha=0.65$ and $\beta=0.34$) from previous work in complex STFT diffusion [27, 28, 4], which scales the amplitudes in the STFT to roughly [1,1][-1,1] while boosting the high frequencies. We define ${\operatorname{T}}(x)={\operatorname{Amp}}(\operatorname{STFT}(x))$ as the transform which maps from waveform to CM input space.
Decoder: $\operatorname{Dec}_{\theta}$ is composed of a U-Net-based denoiser $f_{\theta}$} and the inverse transform ${\operatorname{T}}^{-1}({\mathbf{x}})$. The denoiser reconstructs ${\mathbf{x}}={\operatorname{T}}(x)$ from a noise-corrupted version ${\mathbf{x}}_{\sigma}={\mathbf{x}}+\sigma\bm{\epsilon}$ ($\bm{\epsilon}\sim\mathcal{N}(0,I)$), conditioned on the latent ${\mathbf{z}}_{x}$ and the noise level $\sigma$. ${\mathbf{z}}_{x}$ is upsampled with a dedicated network mirroring the U-Net’s upsampling block and added to each layer of the U-Net. Noise level information is added to every layer via positional embeddings. $f_{\theta}$} is parameterized by a noise prediction network $F_{\theta}$} with a skip connection weighted by coefficients ($c_{\text{skip}},c_{\text{out}}$) that depend on $\sigma$ [29], which additionally enforces the boundary condition necessary for CT [17, 18, 4]:
@@ 3.2 Learning Linearity in CAEs via Implicit Regularization
Implicit homogeneity: Inspired by [1], we apply a random positive gain ($\color[rgb]{1,0,1}a$}}) to the latent and task the decoder to reconstruct a scaled version of the input (${{\color[rgb]{1,0,1}a}}x$). The adapted CT training, depicted in Figure 2(2(b)), uses two parallel pathways. $\operatorname{Enc}_{\theta}$ encodes unscaled $x$ to obtain ${\mathbf{z}}_{x}$. Then, the denoiser ($f_{\theta}$}) receives a noisy scaled input (${\operatorname{T}}({{\color[rgb]{1,0,1}a}}x)+\sigma\bm{\epsilon}$) conditioned on the scaled latent (${{\color[rgb]{1,0,1}a}}{\mathbf{z}}_{x}$). $\color[rgb]{1,0,1}a$}} is never provided as an explicit input to any part of the model, so $\operatorname{Dec}_{\theta}$ must learn to infer the correct output scale from the magnitude of the conditioning latent. If ${{\color[rgb]{1,0,1}a}}=1$, we recover the original CAE training.
Implicit additivity: Inspired by Mixit [30], we augment our data by creating artificial mixes from pairs of elements $(u,v)$ randomly selected from the training set. When decoding, we replace the true latent of the artificial mix ($\operatorname{Enc}_{\theta}(u+v)$) with the sum of the latents of the individual signals: ${\color[rgb]{0,0,1}{\mathbf{z}}^{\prime}}\leftarrow({\color[rgb]{0,0,1}{\mathbf{z}}_{u}+{\mathbf{z}}_{v}})$, where $({\mathbf{z}}_{u}=\operatorname{Enc}_{\theta}(u),{\mathbf{z}}_{v}=\operatorname{Enc}_{\theta}(v))$, as illustrated in Figure 2(2(c)). The denoiser is thus conditioned on $\color[rgb]{0,0,1}{\mathbf{z}}^{\prime}$} and tasked with denoising ${\operatorname{T}}({{\color[rgb]{1,0,1}a}}(u+v))+\sigma\bm{\epsilon})$. We create mixtures from each training batch by concatenating it with a version of itself that has been circularly shifted by one position and summed with the original (Figure 2(2(d))).
@@ 3.3 Training
With the modifications described in Section 3.2 highlighted in color, we train the model under the CT objective $\mathcal{L}_{\text{CT}}$ [17, 18, 4], which minimizes the distance between the denoised output of a teacher ($f_{\theta^{-}}$) and student ($f_{\theta}$) models ($\theta^{-}\leftarrow\text{stopgrad}(\theta)$) for noisy inputs at two different noise levels, $\sigma_{t_{1}}$ and $\sigma_{t_{2}}$.
$\begin{split}\mathcal{L}_{\text{CT}}=\mathbb{E}_{{\mathbf{x}},\bm{\epsilon},t_{1},{{\color[rgb]{1,0,1}a}}}\bigg[\lambda\cdot d\Big(f_{\theta}&({\operatorname{T}}({{\color[rgb]{1,0,1}a}}x)+\sigma_{t_{2}}\bm{\epsilon},\sigma_{t_{2}},{{\color[rgb]{1,0,1}a}}{{\color[rgb]{0,0,1}{\mathbf{z}}}}^{\prime}),\\ f_{\theta^{-}}&({\operatorname{T}}({{\color[rgb]{1,0,1}a}}x)+\sigma_{t_{1}}\bm{\epsilon},\sigma_{t_{1}},{{\color[rgb]{1,0,1}a}}{{\color[rgb]{0,0,1}{\mathbf{z}}}}^{\prime})\Big)\bigg],\end{split}$ (4)
where $\color[rgb]{0,0,1}{\mathbf{z}}$$^{\prime}$ is either ${\mathbf{z}}_{x}$ or ${\mathbf{z}}_{u}+{\mathbf{z}}_{v}$ depending on whether the input waveform is sampled from the training set or is an artificial mixture ($x=u+v$), $\bm{\epsilon}\sim\mathcal{N}(0,I)$ is a shared noise direction, $\lambda(\sigma_{t_{1}},\sigma_{t_{2}})=\frac{1}{\sigma_{t_{2}}-\sigma_{t_{1}}}$ is a weighting function designed to give higher weights to smaller noise increments [18]. We sample time steps $t_{1}\in[0,1]$ and corresponding noise levels $\sigma_{t_{1},t_{2}}$ following [4], with $t_{2}=t_{1}+\Delta_{t_{k}}$, where $\Delta_{t_{k}}$ depends on the training step $k$ and decays over training ($\Delta t_{k}=\Delta t_{0}^{\frac{k}{K}(e_{K}-1)+1}$, $\Delta t_{0}=0.1$, $e_{K}=2.0$). $d$ is the Pseudo-Huber loss [31]: $d(x,y)=\sqrt{|x-y|^{2}+c^{2}}-c$, with c=5.4e4c=5.4e-4.
@@ 4.1 Experimental setup
Data: The training corpus is a large-scale music/speech dataset compiled from MTG-Jamendo [32], MoisesDB [33], M4Singer[34], DNS-Challenge [35], and E-GMD [36]. Tracks are sampled with weights $(60,20,9,8,3)$ respectively, and a segment of $2$ seconds is randomly cropped from each track. Segments are converted to mono and resampled to $44.1$ kHz.
Baselines: We report metrics using the publicly available weights of Music2Latent (M2L-Pub) [4] . As a fairer baseline, we retrain M2L-Pub on the same data as our model, including the random gains and artificial mixtures for data augmentation, but without our proposed implicit homogeneity and additivity conditioning strategies (M2L). We also report metrics for the public weights from Stable Audio $1.0$ VAE [11] (SA-VAE), serving as another autoencoder baseline with a different architecture and training procedure.
@@
M2L-Pub [4] $5.01$ 0.79-0.79 $2.82$ $1.16$ 1.29-1.29 $95.34$ 14.85-14.85 $28.08$ 16.30-16.30 $6.78$ 16.11-16.11 $5.10$ 18.16-18.16
SA-VAE [11] $5.38$ 12.58-12.58 $1.71$ $\mathbf{0.57}$ $\mathbf{4.03}$ $323.26$ $\underline{-2.81}$ $87.80$ $\underline{-3.11}$ $3.65$ $\underline{-4.10}$ $3.08$ $\underline{-4.48}$
@@
M2L-Pub [4] $1.14$ $1.85$ $2.52$ 4.69-4.69 $12.13$ $\mathbf{5.69}$
SA-VAE [11] $\mathbf{0.72}$ $\mathbf{7.32}$ $3.03$ $\underline{-1.27}$ $\underline{4.59}$ $6.27$

### v2

@@ 1 Introduction
Modern Autoencoders can achieve excellent reconstruction quality at high compression rates at the expense of complex, entangled latent spaces. For applications where input space manipulation is desirable from within the compressed space, recent research has proposed either task-specific adaptors for pre-trained models or redesigning the autoencoder to preserve key structural properties, such as equivariance to spatial transformations [kouzelisEQVAEEquivarianceRegularized2025, skorokhodovImprovingDiffusabilityAutoencoders2025, zhouAliasfreeLatentDiffusion2025]. This work follows the latter approach.
@@
This work presents a training methodology for constructing an approximately linear compressed audio representation that enables intuitive manipulation, where simple algebraic operations in the latent space correspond directly to mixing and scaling in the audio domain. The method employs implicit regularization through data augmentation, without modifying the model architecture or objective. The approach is demonstrated with the Music2Latent architecture [pasiniMusic2LatentConsistencyAutoencoders2024], a Consistency Autoencoder (CAE) that achieves high-quality, single-step reconstruction and a $64\times$ compression rate for $44.1$ kHz audio.
@@ 2.1 Audio processing in the latent space
There is a growing interest in manipulating audio by training task-specific modules which operate directly in the latent space of a pretrained AE, with applications in source separation [mancusiUnsupervisedSourceSeparation2021, postolacheLatentAutoregressiveSource2023, bindiUnsupervisedComposableRepresentations2024], speech enhancement [omranDisentanglingSpeechSurroundings2023, liSpeechEnhancementUsing2025], upsampling and upmixing [braliosLearningUpsampleUpmix2025], filtering [braliosRebottleneckLatentRestructuring2025], and generative modelling [schneiderMousaiEfficientTexttomusic2024, evansStableAudioOpen2025, nistalDiffariffMusicalAccompaniment2024]. Some works retrain AEs for specific objectives, such as source disentanglement [bieLearningSourceDisentanglement2025] or separation via latent masking [tzinisTwostepSoundSource2020]. In contrast, our work focuses on training an AE to have desirable, task-agnostic structural properties. This enables direct and efficient audio manipulation and can serve as a strong foundation for these downstream applications.
@@ 2.2 Diffusion and Consistency models
Denoising Diffusion Probabilistic Models [hoDenoisingDiffusionProbabilistic2020a] and score-based models [songScorebasedGenerativeModeling2021] have achieved great success in generative modeling, where the goal is to estimate the underlying data distribution from samples. These models define a forward process that gradually adds noise to data and learn a reverse process to denoise it using a neural network. Sampling typically requires an iterative procedure with many steps to generate a clean sample. Consistency Models [songConsistencyModels2023] accelerate this process by mapping any point along the trajectory defined by the probability flow ordinary differential equation [songScorebasedGenerativeModeling2021] directly to the origin (the clean data point), enabling single-step generation. When trained from scratch, the process is called Consistency Training (CT), in which a ”student” denoiser network ($f_{\theta}$) is trained to match the output of a ”teacher” ($f_{\theta^{-}}$), which itself denoises a less corrupted version of the same data point. In improved Consistency Training (iCT) [songImprovedTechniquesTraining2023], the teacher is updated with the same parameters $\theta$ as the student but detached from the computational graph ($\theta^{-}\leftarrow\text{stopgrad}(\theta)$).
@@ 2.3 High-Fidelity Audio Autoencoders
In audio, popular AEs include Neural Audio Codecs [defossezHighFidelityNeural2023], which compress audio into a set of discrete tokens, and Variational Autoencoders [kingmaAutoencodingVariationalBayes2014], which have seen significant success in generative modeling. However, both frameworks require complex, multi-stage training with adversarial objectives [esserTamingTransformersHighresolution2021] to achieve high-quality reconstructions. Diffusion Autoencoders [preechakulDiffusionAutoencodersMeaningful2022, birodkarSampleWhatYou2024, chenDiffusionAutoencodersAre2025, zhaoEpsilonvaeDenoisingVisual2024, schneiderMousaiEfficientTexttomusic2024] enable high-fidelity reconstruction by replacing the deterministic decoder with a conditional diffusion model. Notable examples in audio include [schneiderMousaiEfficientTexttomusic2024] and Music2Latent [pasiniMusic2LatentConsistencyAutoencoders2024, pasiniMusic2Latent2AudioCompression2025]. A specific instance of DiffAEs is the CAE[pasiniMusic2LatentConsistencyAutoencoders2024, pasiniMusic2Latent2AudioCompression2025], where the decoder is a CM, enabling decoding in one step. DiffAEs can be trained with a single diffusion/CT objective [preechakulDiffusionAutoencodersMeaningful2022, chenDiffusionAutoencodersAre2025, pasiniMusic2LatentConsistencyAutoencoders2024]. High-quality decoding has been linked to the capacity of sampling details at inference-time, since the decoder acts as a denoiser instead of an upsampler, which also reduces the amount of unnecessary information to be encoded in the latent [birodkarSampleWhatYou2024, zhaoEpsilonvaeDenoisingVisual2024].
@@ 3 Method
We postulate that approximate linearity can be achieved through a simple data augmentation scheme, without the need to modify the model or loss function. The intuition is that by performing latent space operations (scaling and addition) and asking the decoder to recover the corresponding audio signal, it learns to behave linearly. Our approach is, in theory, model-agnostic and can be applied to any AE. In this work, we apply it to a CAE architecture for audio [pasiniMusic2LatentConsistencyAutoencoders2024], which offers $64\times$ compression and high-quality reconstruction with a single training objective.
@@ 3.1 Model architecture
Our model (Lin-CAE) follows the Music2Latent architecture [pasiniMusic2LatentConsistencyAutoencoders2024], illustrated in Figure 2(2(a)). We recall the main components here for clarity, but we refer the reader to [pasiniMusic2LatentConsistencyAutoencoders2024] for the full details.
CAE Denoiser representation space: We use a complex-valued Short-Time Fourier Transform (STFT) followed by the invertible amplitude scaling ${\operatorname{Amp}}:{\mathbb{C}}\to{\mathbb{C}}=\beta|c|^{\alpha}e^{i\angle(c)}$ (with $\alpha=0.65$ and $\beta=0.34$) from previous work in complex STFT diffusion [richterSpeechEnhancementDereverberation2023, zhuEdmsoundSpectrogramBased2023, pasiniMusic2LatentConsistencyAutoencoders2024], which scales the amplitudes in the STFT to roughly [1,1][-1,1] while boosting the high frequencies. We define ${\operatorname{T}}(x)={\operatorname{Amp}}(\operatorname{STFT}(x))$ as the transform which maps from waveform to CM input space.
Decoder: $\operatorname{Dec}_{\theta}$ is composed of a U-Net-based denoiser $f_{\theta}$} and the inverse transform ${\operatorname{T}}^{-1}({\mathbf{x}})$. The denoiser reconstructs ${\mathbf{x}}={\operatorname{T}}(x)$ from a noise-corrupted version ${\mathbf{x}}_{\sigma}={\mathbf{x}}+\sigma\cdot\bm{\epsilon}$ ($\bm{\epsilon}\sim\mathcal{N}(0,I)$), conditioned on the latent ${\mathbf{z}}_{x}$ and the noise level $\sigma$. ${\mathbf{z}}_{x}$ is upsampled with a dedicated network mirroring the U-Net’s upsampling block and added to each layer of the U-Net. Noise level information is added to every layer via positional embeddings. $f_{\theta}$} is parameterized by a noise prediction network $F_{\theta}$} with a skip connection weighted by coefficients ($c_{\text{skip}},c_{\text{out}}$) that depend on $\sigma$ [karrasElucidatingDesignSpace2022], which additionally enforces the boundary condition necessary for CT [songConsistencyModels2023, songImprovedTechniquesTraining2023, pasiniMusic2LatentConsistencyAutoencoders2024]:
@@ 3.2 Learning Linearity in CAEs via Implicit Regularization
Implicit homogeneity: Inspired by [kouzelisEQVAEEquivarianceRegularized2025], we apply a random positive gain ($\color[rgb]{1,0,1}a$}}) to the latent and task the decoder to denoise a scaled version of the input (${{\color[rgb]{1,0,1}a}}\cdot x$). The adapted CT training, depicted in Figure 2(2(b)), uses two parallel pathways. $\operatorname{Enc}_{\theta}$ encodes unscaled $x$ to obtain ${\mathbf{z}}_{x}$. Then, the denoiser ($f_{\theta}$}) receives a noisy scaled input (${\operatorname{T}}({{\color[rgb]{1,0,1}a}}\cdot x)+\sigma\cdot\bm{\epsilon}$) conditioned on the scaled latent (${{\color[rgb]{1,0,1}a}}\cdot{\mathbf{z}}_{x}$). $\color[rgb]{1,0,1}a$}} is never provided as an explicit input to any part of the model, so $\operatorname{Dec}_{\theta}$ must learn to infer the correct output scale from the magnitude of the conditioning latent. If ${{\color[rgb]{1,0,1}a}}=1$, we recover the original CAE training.
Implicit additivity: Inspired by Mixit [wisdomUnsupervisedSoundSeparation2020], we augment our data by creating artificial mixes from pairs of elements $(u,v)$ randomly selected from the training set. When decoding, we replace the true latent of the artificial mix ($\operatorname{Enc}_{\theta}(u+v)$) with the average of the latents of the individual signals: ${\color[rgb]{0,0,1}{\mathbf{z}}^{\prime}}\leftarrow 0.5\cdot({\mathbf{z}}_{u}+{\mathbf{z}}_{v})$, where $({\mathbf{z}}_{u}=\operatorname{Enc}_{\theta}(u),{\mathbf{z}}_{v}=\operatorname{Enc}_{\theta}(v))$, as illustrated in Figure 2(2(c)). The denoiser is thus conditioned on $\color[rgb]{0,0,1}{\mathbf{z}}^{\prime}$} and tasked with denoising ${\operatorname{T}}({{\color[rgb]{1,0,1}a}}\cdot 0.5\cdot(u+v))+\sigma\cdot\bm{\epsilon})$. We augment the training batch with mixtures by concatenating it with a version of itself that has been circularly shifted by one position and averaged with the original (Figure 2(2(d))).
@@ 3.3 Training
With the modifications described in Section 3.2 highlighted in color, we train the model under the CT objective $\mathcal{L}_{\text{CT}}$ [songConsistencyModels2023, songImprovedTechniquesTraining2023, pasiniMusic2LatentConsistencyAutoencoders2024], which minimizes the distance between the denoised output of a teacher ($f_{\theta^{-}}$) and student ($f_{\theta}$) models ($\theta^{-}\leftarrow\text{stopgrad}(\theta)$) for noisy inputs at two different noise levels, $\sigma_{t_{1}}$ and $\sigma_{t_{2}}$.
$\begin{split}\mathcal{L}_{\text{CT}}=\mathbb{E}_{{\mathbf{x}},\bm{\epsilon},t_{2},{{\color[rgb]{1,0,1}a}}}\bigg[\lambda\cdot d\Big(f_{\theta}&({\operatorname{T}}({{\color[rgb]{1,0,1}a}}\cdot x)+\sigma_{t_{2}}\cdot\bm{\epsilon},\sigma_{t_{2}},{{\color[rgb]{1,0,1}a}}\cdot{{\color[rgb]{0,0,1}{\mathbf{z}}}}^{\prime}),\\ f_{\theta^{-}}&({\operatorname{T}}({{\color[rgb]{1,0,1}a}}\cdot x)+\sigma_{t_{1}}\cdot\bm{\epsilon},\sigma_{t_{1}},{{\color[rgb]{1,0,1}a}}\cdot{{\color[rgb]{0,0,1}{\mathbf{z}}}}^{\prime})\Big)\bigg],\end{split}$ (4)
where $\color[rgb]{0,0,1}{\mathbf{z}}$$^{\prime}$ is either ${\mathbf{z}}_{x}$ or $0.5\cdot({\mathbf{z}}_{u}+{\mathbf{z}}_{v})$ depending on whether the input waveform is sampled from the training set or is an artificial mixture ($x=0.5\cdot(u+v$)), $\bm{\epsilon}\sim\mathcal{N}(0,I)$ is a shared noise direction, $\lambda(\sigma_{t_{1}},\sigma_{t_{2}})=\frac{1}{\sigma_{t_{2}}-\sigma_{t_{1}}}$ is a weighting function designed to give higher weights to smaller noise increments [songImprovedTechniquesTraining2023]. We sample time steps $t_{2}\in[0,1]$ and corresponding noise levels $\sigma_{t_{1},t_{2}}$ following [pasiniMusic2LatentConsistencyAutoencoders2024], with $t_{1}=\max(t_{2}-\Delta_{t_{k}},0)$, where $\Delta_{t_{k}}$ depends on the training step $k$ and decays over training ($\Delta t_{k}=\Delta t_{0}^{\frac{k}{K}(e_{K}-1)+1}$, $\Delta t_{0}=0.1$, $e_{K}=2.0$). $d$ is the Pseudo-Huber loss [charbonnierDeterministicEdgepreservingRegularization1997]: $d(x,y)=\sqrt{|x-y|^{2}+c^{2}}-c$, with c=5.4e4c=5.4e-4.
@@ 4.1 Experimental setup
Data: The training corpus is a large-scale music/speech dataset compiled from MTG-Jamendo [bogdanovMTGjamendoDatasetAutomatic2019], MoisesDB [pereiraMoisesDBDatasetSource2023], M4Singer[zhangM4SingerMultistyleMultisinger2022], DNS-Challenge [dubeyICASSP2023Deep2023], and E-GMD [callenderImprovingPerceptualQuality2020]. Tracks are sampled with weights $(60,20,9,8,3)$ respectively, and a segment of $2$ seconds is randomly cropped from each track. Segments are converted to mono by averaging stereo channels and resampled to $44.1$ kHz.
Baselines: We report metrics using the publicly available weights of Music2Latent (M2L-Pub) [pasiniMusic2LatentConsistencyAutoencoders2024] . As a fairer baseline, we retrain M2L-Pub on the same data as our model, including the random gains and artificial mixtures for data augmentation, but without our proposed implicit homogeneity and additivity conditioning strategies (M2L). We also report metrics for the public weights from Stable Audio $1.0$ VAE [evansStableAudioOpen2025] (SA-VAE), serving as another autoencoder baseline with a different architecture and training procedure.
@@
M2L-Pub [pasiniMusic2LatentConsistencyAutoencoders2024] $5.01$ 0.79-0.79 $2.82$ $1.16$ 1.29-1.29 $95.34$ 14.85-14.85 $28.08$ 16.30-16.30 $6.78$ 16.11-16.11 $5.10$ 18.16-18.16
SA-VAE [evansStableAudioOpen2025] $5.38$ 12.58-12.58 $1.71$ $\mathbf{0.57}$ $\mathbf{4.03}$ $323.26$ $\underline{-2.81}$ $87.80$ $\underline{-3.11}$ $3.65$ $\underline{-4.10}$ $3.08$ $\underline{-4.48}$
@@
M2L-Pub [pasiniMusic2LatentConsistencyAutoencoders2024] $1.14$ $1.85$ $2.52$ 4.69-4.69 $12.13$ $\mathbf{5.69}$
SA-VAE [evansStableAudioOpen2025] $\mathbf{0.72}$ $\mathbf{7.32}$ $3.03$ $\underline{-1.27}$ $\underline{4.59}$ $6.27$